### v1

@@ 2.3 Data acquisition and reduction - South
All-Stokes photopolarimetry of the Southern dark patches was performed during October of 2021 and August of 2022 on the 1.9-m telescope of the SAAO using HIPPO (Potter et al., 2010). HIPPO’s waveplates are contrarotated at 10 Hz and therefore modulate the ordinary and extraordinary beams through the Thompson beamsplitter. The modulation is sufficiently rapid that errors which arise as a result of variable atmospheric conditions or telescope guiding modulations, become minimal. The systematic uncertainty of is estimated from repeating and comparing observations of standards over multiple seasons. The modulated signal is sampled and recorded by two RCA31034A GaAs photomultiplier tubes every millisecond. See Potter et al. (2010) for more details. All measurements were performed with an R filter in the Kron–Cousin system. Polarized and unpolarized standard stars were observed in order to calculate the position angle offsets, instrumental polarization and efficiency factors. Background sky polarization measurements were also taken at frequent intervals during the observations. Data reduction then proceeded as outlined in Potter et al. (2010).
@@ 3 Results
All measurements of individual targets are corrected for the instrumental error as
and similarly for . The uncertainties of the individual targets include the contribution from instrumental error uncertainty. The final uncertainty is calculated as
and similarly for , where the superscripts ”measured” and ”inst” refer to values obtained from differential photometry and the instrumental error respectively. The instrumental error and its uncertainty is not necessarily identical for Stokes and , nor between targets, as it can be time-dependent. The measurements of each targets were corrected using the values of the instrumental error corresponding to the same time period that the targets were observed.
We present the results of the dark patch candidates in Table 1. We present the measurements for DPN1 in Fig. 1, together with their weighted mean, standard deviation, and error on the weighted mean. We also present as a function of distance in the same figure. Distance information was retrieved from the Gaia DR3 catalog (Gaia Collaboration et al., 2016, 2023), by inverting the provided parallaxes. If parallax information was not available, we used the distance mentioned in the distance_gspphot column of the catalog. Stars with no distance information in either form were not plotted in the figure. Similar plots for all of the fields can be found in Appendix A. The sky location of the dark patches are presented in Fig. 2. Finder charts of the individual stars for which we measured polarization in each region are presented in Appendix B. For each field, we divide the range of RAs and DECs of our targets in four equal bins. For each bin we calculate and plot the average polarization , and its standard deviation. For all of the regions, there is no significant deviation in polarization in any of the bins. Hence, there doesn’t exist spatial dependence of the polarization within the fields, and can be considered constant in polarization. Sky maps were obtained from DSS.
Individual stellar measurements of have average statistical uncertainties of and for the northern fields, and and for the southern. However, the values that characterize the average polarization of each field is the weighted mean of the measurements and the error on the weighted mean, and . Therefore, for each field we calculate the weighted mean of the Stokes parameters as
being the individual measurement of each star and the total number of stars in each field, with , and the uncertainty of the measurement. We are also interested in the weighted standard deviation, defined as
and the error of the weighted mean, which, in the case of measurements with different uncertainties, as ours, is
@@
Polarization is a positively defined quantity, expressed as
with and being the uncertainties of and respectively. Therefore, measurements are biased towards higher values, especially for low signal-to-noise (S/N) measurements (e.g. Vaillancourt, 2006). We debias our measurements using the modified asymptotic (MAS) estimator proposed in Plaszczynski et al. (2014),
with and as defined in Eq. 6. We present the debiased values in Table 1. Conversely, the bias in the polarization angle is likely very small (Montier et al., 2015). Thus the expression of the polarization angle was used without correction.
The observed polarization of stars in any given field is expected to arise solely due to dichroic absorption of the light by aligned dust grains as discussed in Sect. 1. This is based on the fact that most stars are intrinsically unpolarized with the possible exceptions of magnetic stars, or stars surrounded by dusty disks (Fadeyev, 2007; Clarke, 2010). The potential existence of such stars in our samples would manifest as clear outliers in the plots. For each of our patches, we performed a clipping three times, to discard outliers. This process excluded only 1 star from DPN3, and 1 star from DPS7, while the number of the remaining stars per patch ranges between 16 and 37.
The observed scatter between the measurements of different stars in the same region is a combination of the intrinsic scatter of the sources within the region and the statistical scatter introduced by the instrument, arising from photon noise and instrumental systematics. That is
and similar for , with the assumption that these sources of variance are Gaussian. The statistical scatter of both the northern and the southern instruments used for this study is of the order of , as deduced by standards observations. Therefore, the intrinsic scatter in the polarization of the observed patches is maximum for all cases.
Notes. and correspond to the debiased estimate of the mean polarization, and its error respectively. and are similar for the polarization angle. , , and are the weighted mean of values, its standard deviation, and error on the mean respectively. Similar for . Empty cells correspond to information that is not available.
@@ 4 Discussion
We have identified and surveyed twelve regions of the sky, with very low extinction values, to assess whether they exhibit significantly low polarization, to be used as calibrators for wide-field polarimeters. We measured the linear polarization of stars in four patches in the northern, and eight in the southern hemisphere. We find that two regions in the North and seven in the South display mean polarization of , while one more in each hemisphere has a mean polarization marginally above . Moreover, we have included the regions probed by Skalidis et al. (2018) in Table 1 as one of those regions could also serve as a zero-polarization calibrator.
@@
At high Galactic latitudes (as are the regions in our study) the full column density, as well as the dust polarization in emission and absorption, is dominated by the wall of the Local Bubble located between 150 and 300 from the Sun (Skalidis & Pelgrims, 2019; Pelgrims et al., 2020). Most of the observed stars in our samples have distances larger than that. Therefore, we are confident that our stellar polarization measurements sample all the dust column that could induce polarization along the line of sight of the regions surveyed (see plots in Appendix A). Moreover, we do not notice any significant change of the measured polarization as a function of distance. In conclusion, our measurements are characteristic of each field, regardless of distance, and represent the true polarization value arising due to the effect of interstellar dust in each region.
@@
An interesting outcome of this study is the mean polarization of DPN1. Based on its reddening value from Lenz et al. (2017) mag, and the relation , we would not expect its polarization to be at the measured value . This could be due to reasons related to how the dust maps are constructed, and their uncertianties. There is quite high uncertainty in modelling that is associated with dust maps at high latitude. For example, Lenz et al. (2017) compared the Schlegel et al. (1998) maps with data (HI4PI Collaboration et al., 2016) to derive the reddening . This technique of modelling involves uncertainties relating to two factors: i) The gas to dust ratio, which may exhibit variations across the sky (Shull & Panopoulou, 2023; Skalidis et al., 2023) and ii) the dust modeling, which results in important differences at high latitudes (Chiang & Ménard, 2019; Chiang, 2023). Characteristic evidence for the high uncertainties in dust maps, is the differences between maps, especially in high latitudes. For instance, while the reddening value for DPN1 from Lenz et al. (2017) is mag, the corresponding value derived from the Planck Collaboration et al. (2016) map is mag. At the same time, the stellar reddening based maps (e.g. Green et al., 2019) have statistical uncertainties that are comparable to the extinction, so they too cannot be trusted on a pixel-by-pixel basis. Having discussed that, it is worth mentioning that the recent work of Angarita et al. (2023) place higher constrains in the relation , finding an upper limit of in intermediate latititudes . Therefore, the extraordinary value in the mean polarization of DPN1 seems to be mostly associated with the uncertainty of the extinction of the field, but it is not impossible that it is an outlier in the relation.
@@
In conclusion, the most important aspect of this work is that all fields, even the ones with , display very low intrinsic scatter in polarization (). Thus, they can be considered as polarimetric standard fields. Depending on the needs and aims in accuracy of different polarization surveys and instruments, we expect that between 9 and all of the surveyed regions could be used as zero-polarization calibrators for wide-field polarimeters.
All data will become publicly available on CDS.
@@ Acknowledgements.
- Shull & Panopoulou (2023) Shull, J. M. & Panopoulou, G. V. 2023, arXiv e-prints, arXiv:2310.12205
@@ Appendix A Polarization measurements of the observed fields
We present here our measurements of all of the dark patches, similarly to Fig. 1.
@@ Appendix B Finder charts of stars in the observed fields
In this appendix we present the sky maps of the dark patches, where we mark the observed stars. We also plot the dependence of the average polarization between different slices of RA and DEC of the same field, to examine the presence of any spatial dependence of the polarization within the field.

### v2

@@ 2.3 Data acquisition and reduction: South
All-Stokes photopolarimetry of the southern dark patches was performed during October of 2021 and August of 2022 on the 1.9-m telescope of the SAAO, using HIPPO (Potter et al., 2010). HIPPO’s waveplates are contrarotated at 10 Hz and they therefore modulate the ordinary and extraordinary beams through the Thompson beamsplitter. The modulation is sufficiently rapid that errors arising as a result of variable atmospheric conditions or telescope guiding modulations become minimal. The systematic uncertainty of was estimated from repeating and comparing observations of standards over multiple seasons. The modulated signal is sampled and recorded by two RCA31034A GaAs photomultiplier tubes every millisecond. We refer to Potter et al. (2010) for more details. All measurements were performed with an R filter in the Kron–Cousin system. Polarized and unpolarized standard stars were observed to calculate the position angle offsets, instrumental polarization, and efficiency factors. Background sky polarization measurements were also taken at frequent intervals during the observations. The data reduction then proceeded according to the method outlined in Potter et al. (2010).
@@ 3 Results
All measurements of individual targets are corrected for the instrumental error as:
and similarly for . The uncertainties of the individual targets include the contribution from instrumental error uncertainty. The final uncertainty is calculated as:
and similarly for , where the superscripts ”measured” and ”inst” refer to values obtained from differential photometry and the instrumental error, respectively. The instrumental error and its uncertainty are not necessarily identical for Stokes and , nor between targets, as it can be time-dependent. The measurements of each targets were corrected using the values of the instrumental error corresponding to the same time period that the targets were observed.
We present the results of the dark patch candidates in Table 4. We present the measurements for DPN1 in Fig. 1, together with their weighted mean, standard deviation, and error on the weighted mean. We also present as a function of distance in the same figure. Distance information was retrieved from the Gaia DR3 catalog (Gaia Collaboration et al., 2016, 2023), by inverting the provided parallaxes. If parallax information was not available, we used the distance mentioned in the distance_gspphot column of the catalog. Stars with no distance information in either form were not plotted in the figure. Similar plots for all of the fields can be found in Appendix A. The sky location of the dark patches are presented in Fig. 2. Finder charts of the individual stars for which we measured polarization in each region are presented in Appendix B. For each field, we divided the range of RAs and Decs. of our targets in four equal bins. For each bin, we calculated and plotted the average polarization , and its standard deviation. For all of the regions, there is no significant deviation in the polarization in any of the bins. Therefore, there is no spatial dependence of the polarization within the fields and it can be considered constant in terms of polarization. The sky maps were obtained from DSS.
Individual stellar measurements of have average statistical uncertainties of and for the northern fields, and and for the southern. However, the values that characterize the average polarization of each field is the weighted mean of the measurements and the error on the weighted mean, and . Therefore, for each field, we calculate the weighted mean of the Stokes parameters as:
with being the individual measurement of each star and the total number of stars in each field, with , and the uncertainty of the measurement. We are also interested in the weighted standard deviation, defined as:
and the error of the weighted mean, which, in the case of measurements with different uncertainties (as ours) is:
@@
Polarization is a positively defined quantity, expressed as:
with and being the uncertainties of and respectively. Therefore, measurements are biased towards higher values, especially for low signal-to-noise (S/N) measurements (e.g., Vaillancourt, 2006). We debiased our measurements using the modified asymptotic (MAS) estimator proposed in Plaszczynski et al. (2014):
with and defined in Eq. 6. We present the debiased values in Table 4. Conversely, the bias in the polarization angle is likely very small (Montier et al., 2015). Thus, the expression of the polarization angle was used without correction.
The observed polarization of stars in any given field is expected to arise solely due to the dichroic absorption of the light by aligned dust grains (as discussed in Sect. 1). This is based on the fact that most stars are intrinsically unpolarized with the possible exceptions of magnetic stars or stars surrounded by dusty disks (Fadeyev, 2007; Clarke, 2010). The potential existence of such stars in our samples would manifest as clear outliers in the plots. For each of our patches, we performed a clipping three times to discard outliers. This process excluded only 1 star from DPN3, 1 from DPN4, and 1 star from DPS7, while the number of the remaining stars per patch ranges between 16 and 31.
The observed scatter between the measurements of different stars in the same region is a combination of the intrinsic scatter of the sources within the region and the statistical scatter introduced by the instrument, arising from photon noise and instrumental systematics, namely:
and it is similar for , with the assumption that these sources of variance are Gaussian. The statistical scatter of both the northern and the southern instruments used for this study is on the order of , as deduced by standards observations. Therefore, the intrinsic scatter in the polarization of the observed patches is maximum for all cases.
@@ 4 Discussion
We have identified and surveyed twelve regions of the sky with very low extinction values to assess whether they exhibit a sufficiently low polarization to be used as calibrators for wide-field polarimeters. We measured the linear polarization of stars in four patches in the northern hemisphere and eight in the southern. We find that two regions in the north and seven in the south display mean polarization of , while one more in each hemisphere has a mean polarization that is marginally above . Moreover, we have included the regions probed by Skalidis et al. (2018) in Table 4 as one of those regions could also serve as a zero-polarization calibrator.
@@
At high Galactic latitudes (i.e., the regions in our study), the full column density, as well as the dust polarization in emission and absorption are dominated by the wall of the Local Bubble located between 150 and 300 from the Sun (Skalidis & Pelgrims, 2019; Pelgrims et al., 2020). Most of the observed stars in our samples have distances larger than that. Therefore, we are confident that our stellar polarization measurements sample all the dust column that could induce polarization along the LoS of the regions surveyed (see plots in Appendix A). Moreover, we do not notice any significant change of the measured polarization as a function of distance. In conclusion, our measurements are characteristic of each field, regardless of distance, and represent the true polarization value arising due to the effect of interstellar dust in each region.
@@
An interesting outcome of this study is the mean polarization of DPN1. Based on its reddening value from Lenz et al. (2017) mag and the relation , we would not expect its polarization to be at the measured value . This could be due to reasons related to how the dust maps are constructed and their uncertainties. There is quite high uncertainty in the modeling that is associated with dust maps at high latitude. For example, Lenz et al. (2017) compared the Schlegel et al. (1998) maps with data (HI4PI Collaboration et al., 2016) to derive the reddening . This technique of modeling involves uncertainties relating to two factors: i) the gas-to-dust ratio, which may exhibit variations across the sky (Shull & Panopoulou, 2024; Skalidis et al., 2023), and ii) the dust modeling, which results in important differences at high latitudes (Chiang & Ménard, 2019; Chiang, 2023). Characteristic evidence for the high uncertainties in the dust maps is revealed by the differences between maps, especially at high latitudes. For instance, while the reddening value for DPN1 from Lenz et al. (2017) is mag, the corresponding value derived from the Planck Collaboration et al. (2016) map is mag. At the same time, the stellar reddening based maps (e.g., Green et al., 2019) have statistical uncertainties that are comparable to the extinction, so they too cannot be trusted on a pixel-by-pixel basis. Having discussed this peculiarity, it is worth mentioning that the recent work of Angarita et al. (2023) placed higher constraints on the relation , finding an upper limit of in intermediate latititudes . Therefore, the extraordinary value in the mean polarization of DPN1 seems to be mostly associated with the uncertainty of the extinction of the field, but it is not impossible for it to be an outlier in the relation.
@@
In conclusion, the most important aspect of this work is that all fields, even the ones with , display very low intrinsic scatter in polarization (). Thus, they can be considered as polarimetric standard fields. Depending on the needs and aims in accuracy of different polarization surveys and instruments, we expect that a total between 9 and all 12 of the surveyed regions could be used as zero-polarization calibrators for wide-field polarimeters. All the data in this study are to be made publicly available on CDS.
@@ Acknowledgements.
- Shull & Panopoulou (2024) Shull, J. M. & Panopoulou, G. V. 2024, ApJ, 961, 204
@@ Appendix A Polarization measurements of the observed fields
We present our measurements of all of the dark patches, similarly to Fig. 1.
@@ Appendix B Finder charts of stars in the observed fields
In this appendix, we present the sky maps of the dark patches, where we mark the observed stars. We also plot the dependence of the average polarization between different slices of RA and Dec of the same field to examine the presence of any spatial dependence of the polarization within the field.